A 95% confidence interval obtained from a sample of 100 outpatients for the true population mean normal mean systolic blood pressure is given by (114 mmHG, 120 mmHG).

A) Find 99% Confidence Interval

B) Can you tell how much would it be wider compare to 95% Confidence Interval?

Mention all the steps for the both answers.
Answer #1

l14, 120 Substitute tha nluyin 文+E= 120 120-11-7二3 Value Value Coyves lin lVeIO存 -) 117 +0.96) ( 12.0 = loo σ (120-117 0.196 3 0-196 σ= 15-306) A) we knouo Fron the Strda d noinal table, the c ; 6cal value Comes Suhstibuke tua values n tue alove ormula (5. 306l IO0 /C113-067, 120943(B) Siven that 14120) qac/o înterval.foY -The confidence İntenal fDY 1113,057,120 ,943), 6 Cov the q9f cT, the ciffeie)ca btp